# FuseGPT: Learnable Layers Fusion of Generative Pre-trained Transformers

### v1

@@ IV Experiments
TABLE III: Zero-shot task results for multimodal models. Randomly select samples from WikiText-2 training dataset as calibration data.
@@ IV-A Experimental Setting
FuseGPT is implemented based on Hugging Face Transformers [35] together with Pytorch [36]. The experiments are conducted on one NVIDIA H800 PCIe 80GB graphics card with CUDA Driver 12.6. When deployed for evaluation, we can explicitly compute and store $\mathbf{W}=\mathbf{W}+\mathbf{C}\odot\mathbf{W}_{p}$ and perform inference as usual, thus no additional costs are produced. We randomly selected samples from WikiText-2 training dataset [37] as calibration and fine-tuning data. Without special description, we use $32$ samples for calibration and $1024$ samples for fine-tuning, which is extremely lightweight for model compression. We use the partial group size $G$ as $7$, which enables us to only update approximately $25\%$ parameters for a 7B model. We set the rank for decomposition of coefficient $\mathbf{C}$ as $128$. In order to further reduce learning costs, we also employ LoRA [33] with rank $128$ to update the origin weights inside the partial group. We run $20$ epochs during the update of partial group parameters. We use the Adam optimizer [38] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.95 and the cosine learning rate decay learning rate scheduler [39]. Specifically, we set different initial learning rate for the decomposed coefficients and other parameters, i.e. $0.001$ and 9.65e-6, respectively. The batch size for partial group fine-tuning is $8$. Based on the above setting, it takes approximately $1.5$ hours for the fusion of LLaMA-2-7B with $25\%$ sparsity.
IV-B Main Results
Generation Performance
@@ IV-B Main Results
We evaluate the perplexity of the pruned models by various sparsity levels. If the product of the total number of transformer blocks in a model and the target sparsity is not an integer, we round up the result to determine the number of transformer blocks to remove as in previous works. As the choice of the calibration data can have a significant influence on the perplexity results [14], we demonstrate the results on two different datasets, i.e. WikiText-2 [37] and C4 [40]. We evaluate the performance of the LLaMA model families across different sizes, including LLaMA-2 [3] and LLaMA-3 [41], and additionally evaluate LLaVA-1.5 models [42]. FuseGPT is compared with various baselines, including ShortGPT [12], SliceGPT [27] and SLEB [14].
@@
Table I shows the perplexity obtained by various sparsity levels. It can be observed that, compared with ShortGPT and SLEB, which are also block pruning techniques, FuseGPT can always achieve lower perplexity. It is important to note that though the sparsity of FuseGPT is 25%, the perplexity is even lower than that in previous works with a sparsity of 20%. Such results demonstrate the state-of-the-art (SOTA) performance on depth pruning of GPTs. When compared with SliceGPT, the technique to reduce the weight size and embedding dimension, FuseGPT also presents competitive performance. It can be observed that SliceGPT works well on the WikiText-2 dataset, but its performance largely decreases for the C4 dataset. However, FuseGPT can not only give competitive results on WikiText-2, but also achieve the SOTA results on C4, which is not inside the fine-tuning dataset. This phenomenon demonstrates the remarkable generalization ability of FuseGPT, which is attributed to the success of recycling the redundant blocks. It should also be noted that FuseGPT can also work well on maintaining the generation performance for large multi-modal models, which proves the wide applicability of FuseGPT.
Zero-shot Experiments
@@
We also evaluate the performance of the pruned models on zero-shot tasks. For language models, we use the following five benchmarks as previous works: PIQA [43]; WinoGrande [44]; HellaSwag [45]; ARC-e and ARC-c [46]. And we also use LM Evaluation Harness [47] with default parameters for evaluations.
@@
Table II shows the results of FuseGPT compared with SLEB and SliceGPT on these tasks. It can be observed that, for LLaMA-2 models, FuseGPT can always achieve better accuracy when the sparsity is set as 25%, with approximately 2% higher accuracy. It is also interesting to find that when FuseGPT prunes the GPTs with a sparsity of 30%, the accuracy gap to the 25% sparsity result of SLEB is less than 1%. For the LLaMA-3 model, FuseGPT can achieve even better results by outperforming previous works by nearly 4%.
@@
For multimodal models, we select five well-known benchmarks: MMMU [48], CMMMU [49], GQA [50], AI2D [51], and OK-VQA [52]. As demonstrated in Table III, we can observe that FuseGPT still achieves state-of-the-art performance. Despite the pruning of 25% parameters, the pruned models can still maintain good performance on various tasks.
@@
IV-C Other Experiments
Impact of Layers Fusion
We conduct an ablation study to show the contribution of layers fusion to the final results. We set up the experiments under three configurations: 1 “MI”: Only keep the importance detection process in FuseGPT; 2 “MI + FT”: Further combine MI with trivial partial group fine-tuning; 3 “MI + Fusion”: The complete FuseGPT that combines importance detection with learnable layers fusion. We use 32 samples for calibration and 256 samples for fine-tuning, which are randomly selected from WikiText-2. We focus on the case of sparsity of 25% and measure the perplexity.
The results in Table IV illustrate that our method to fuse the layers into other blocks indeed helps the performance recovery and it is better than simply updating the parameters. It should be noted that, fine-tuning on the Wiki-Text-2 training set can largely reduce the perplexity when evaluating on it, while it has no negative effects to the performance on C4 dataset. Meanwhile, the results emphasize the importance to choose fine-tuning dataset in real-world scenarios, which can significantly decide the directions of learning and thus the performance.
TABLE IV: Impact of Layers Fusion.
Model Method WikiText-2 C4
LLaMA-2-7B MI 10.35 13.34
MI + FT 7.87 11.33
MI + Fusion 7.44 11.30
LLaVA-1.5-7B MI 11.84 15.06
MI + FT 9.19 13.24
MI + Fusion 8.79 13.22
Figure 3: The Perplexity results on C4 with different dataset size.
Impact of Dataset Size
We conduct an ablation study to show the influence of dataset size on the final results. We set up the experiments by arranging different dataset sizes for calibration and fine-tuning. We choose other two methods for comparison with FuseGPT, and evaluate the perplexity variation on different dataset sizes. We change the calibration dataset size for SliceGPT and SLEB. For FuseGPT, we fix 32 samples for calibration and change the size of the fine-tuning dataset. The samples are randomly selected from WikiText-2, and we focus on the case of sparsity of 25%.
As illustrated in Fig. 3, we can find that the perplexities of the pruned model from FuseGPT are always lower than other methods. It is noteworthy that the perplexities don’t change too much when the dataset size increases. The reason may be that the training data, i.e. WikiText-2, has different distributions from C4, which hinders the performance recovery by adding dataset size. This phenomenon may also indicate that only a small amount of data is sufficient for block importance detection and performance recovery.
V Conclusion
In this paper, we introduced FuseGPT, an innovative approach to structured pruning that enhances the deployment of Generative Pre-trained Transformers (GPTs) in resource-constrained environments. By focusing on the efficient recycling and fusion of redundant transformer blocks, FuseGPT addresses the critical challenge of maintaining high-performance post-pruning. Our methodology leverages a novel metric, Macro Inference (MI), to assess the long-term influence of block removal, ensuring a more informed and effective pruning process. By integrating unimportant blocks into neighboring ones, we preserve valuable pre-trained knowledge, thus minimizing performance degradation. The promising results of FuseGPT highlight its potential as a significant advancement in model compression techniques. It not only facilitates efficient deployment but also opens new avenues for future research in optimizing large-scale GPTs. This work contributes to the ongoing efforts to balance model complexity with practical deployment needs, paving the way for more accessible and sustainable AI solutions.

### v2

@@ 4 Experiments
Table 3: Zero-shot task results for multimodal models. Randomly select samples from WikiText-2 training dataset as calibration data.
@@ 4.1 Experimental Setting
FuseGPT is implemented based on Hugging Face Transformers [46] together with Pytorch [36]. When deployed for evaluation, we can explicitly compute and store $\mathbf{W}=\mathbf{W}+\mathbf{C}\odot\mathbf{W}_{p}$ and perform inference as usual, thus no additional costs are produced. We randomly selected samples from the WikiText-2 training dataset [35] as calibration and fine-tuning data. Without special description, we use $32$ samples for calibration and $1024$ samples for fine-tuning, which is extremely lightweight for model compression. We use the partial group size $G$ as $7$, which enables us to only update approximately $25\%$ parameters for a 7B model. We set the rank for decomposition of coefficient $\mathbf{C}$ as $128$. In order to further reduce learning costs, we also employ LoRA [16] with rank $128$ to update the origin weights inside the partial group. We run $20$ epochs during the update of partial group parameters. We use the Adam optimizer [20] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.95 and the cosine learning rate decay learning rate scheduler [29]. Specifically, we set different initial learning rates for the decomposed coefficients and other parameters, i.e., $0.001$ and 9.65e-6, respectively. The batch size for partial group fine-tuning is $8$. Based on the above setting, it takes approximately $1.5$ hours for the fusion of LLaMA-2-7B with $25\%$ sparsity, where the sparsity is defined as in previous works: (#pruned blocks)/(#blocks in dense model).
4.2 Main Results
@@
Generation Performance. We evaluate the perplexity of the pruned models by various sparsity levels. If the product of the total number of transformer blocks in a model and the target sparsity is not an integer, we round up the result to determine the number of transformer blocks to remove, as in previous works. As the choice of the calibration data can have a significant influence on the perplexity results [41], we demonstrate the results on two different datasets, i.e. WikiText-2 [35] and C4 [38]. We evaluate the performance of the LLaMA model families across different sizes, including LLaMA-2 [43] and LLaMA-3 [6], and additionally evaluate LLaVA-1.5 models [25]. FuseGPT is compared with various baselines, including ShortGPT [34], SliceGPT [1], and SLEB [41].
@@
Table 1 shows the perplexity obtained by various sparsity levels. It can be observed that, compared with ShortGPT and SLEB, which are also block pruning techniques, FuseGPT can always achieve lower perplexity. It is important to note that though the sparsity of FuseGPT is 25%, the perplexity is even lower than that in previous works with a sparsity of 20%. Such results demonstrate the state-of-the-art (SOTA) performance on depth pruning of GPTs. When compared with SliceGPT, the technique to reduce the weight size and embedding dimension, FuseGPT also presents competitive performance. It can be observed that SliceGPT works well on the WikiText-2 dataset, but its performance largely decreases for the C4 dataset. However, FuseGPT can not only give competitive results on WikiText-2 but also achieve the SOTA results on C4, which is not inside the fine-tuning dataset. This phenomenon demonstrates the remarkable generalization ability of FuseGPT, which is attributed to the success of recycling the redundant blocks. It should also be noted that FuseGPT can also work well on maintaining the generation performance for large multi-modal models, which proves the wide applicability of FuseGPT.
@@
Zero-shot Experiments. We also evaluate the performance of the pruned models on zero-shot tasks. For language models, we use the following five benchmarks as previous works: PIQA [2]; WinoGrande [39]; HellaSwag [51]; ARC-e and ARC-c [5]. We also use LM Evaluation Harness [10] with default parameters for evaluations.
@@
Table 2 shows the results of FuseGPT compared with SLEB and SliceGPT on these tasks. It can be observed that, for LLaMA-2 models, FuseGPT can always achieve better accuracy when the sparsity is set as 25%, with approximately 2% higher accuracy. It is also interesting to find that when FuseGPT prunes the GPTs with a sparsity of 30%, the accuracy gap to the 25% sparsity result of SLEB is less than 1%. For the LLaMA-3 model, FuseGPT can achieve even better results by outperforming previous works by nearly 4%.
@@
For multimodal models, we select five well-known benchmarks: MMMU [50], CMMMU [52], GQA [17], AI2D [18], and OK-VQA [33]. As demonstrated in Table 3, we can observe that FuseGPT still achieves state-of-the-art performance. Despite the pruning of 25% parameters, the pruned models can still maintain good performance on various tasks.
@@
4.3 Other Experiments
Performance of MI and Layers Fusion. Table 4 presents three experimental settings, showcasing the generality of MI and the gains from adding layers fusion. The first block compares individual pruning criteria, the second block reflects the same methods combined with LoRA fine-tuning, and the third block evaluates each method integrated with our fusion technique. Comparison of individual criteria. We observe that MI consistently achieves lower perplexities than BI [34], SliceGPT [1], and SLEB [41], even when using fewer calibration samples. For instance, MI at just eight samples yields perplexities comparable to SLEB at 128 samples, which underscores MI’s robustness in identifying redundant blocks under data constraints. LoRA fine-tuning. Augmenting each method with 10-epoch LoRA training lowers the perplexities, yet performance gaps remain. For example, SliceGPT + LoRA shows improvement over pure SliceGPT on WikiText-2 but still underperforms on C4. In contrast, MI + LoRA maintains strong performance across both benchmarks, implying that MI-based pruning retains crucial knowledge for downstream adaptation. Effect of layers fusion. Fusing parameters from pruned blocks leads to substantial perplexity reductions, especially on the more challenging C4 dataset. Notably, SLEB + Fusion outperforms SLEB + LoRA, showing that parameter recycling can exceed the benefits of additional fine-tuning alone. Finally, MI + Fusion consistently gives the best results on WikiText-2 and C4, demonstrating that coupling MI’s informed block selection with layers fusion effectively preserves model capacity and mitigates the degradation typically caused by pruning.
Table 4: Performance of MI and Layers Fusion.
Method # data WikiText-2 C4
BI 128 25.44 31.67
SliceGPT 128 7.56 31.62
SLEB 128 10.39 13.74
MI 8 10.52 13.24
MI 32 10.35 13.34
MI 128 10.26 13.30
BI + LoRA 1024 8.11 16.51
SliceGPT + LoRA 1024 6.32 32.09
SLEB + LoRA 1024 7.48 14.87
MI + LoRA 1024 7.79 15.08
BI + Fusion 1024 12.55 16.08
SLEB + Fusion 1024 7.28 12.40
MI + Fusion 128 7.44 11.30
MI + Fusion 1024 7.19 11.17
Table 5: Comparison with other pruning methods.
Method WikiText-2 C4 latency (ms) Speedup
Dense 5.27 7.27 111.73 -
SparseGPT (2:4) 8.67 14.73 101.57 1.10$\times$
Wanda (2:4) 11.35 16.22 101.57 1.10$\times$
LLM-Pruner (25%) 10.58 12.25 98.87 1.13$\times$
SliceGPT (25%) 7.56 31.62 98.87 1.13$\times$
FuseGPT(25%) 7.19 11.17 84.42 1.33$\times$
Comparison with other pruning methods. Table 5 compares FuseGPT with unstructured (2:4) sparsity approaches (SparseGPT [8] and Wanda [42]) and structured channel-pruning techniques (LLM-Pruner [31] and SliceGPT [1]). All results are measured by running a 2K-token sequence on a single GPU. While SparseGPT and Wanda provide moderate speedups (1.10×) at the cost of increased perplexity, LLM-Pruner and SliceGPT—though structured at the channel dimension—still retain higher perplexities (10.58–12.25 and 7.56–31.62, respectively).
By contrast, FuseGPT (25% block pruning) obtains the lowest perplexity on WikiText-2 (7.19) and C4 (11.17) while achieving a 1.33× speedup over the dense model. These results suggest that block-level pruning combined with proper weight recycling can better preserve model quality than channel-wise strategies while offering greater runtime acceleration than unstructured sparsity methods.
Table 6: FuseGPT + 4-bit weight quantization.
Method # data WikiText-2 C4
FuseGPT 128 7.44 11.30
FuseGPT + GPTQ 128 7.86 11.77
FuseGPT 1024 7.19 11.17
FuseGPT + GPTQ 1024 7.51 11.80
Combined with Quantization. Table 6 reports results of combining FuseGPT with GPTQ-based 4-bit quantization [9]. Despite the drastically reduced precision, FuseGPT + GPTQ suffers only a modest perplexity increase relative to the unquantized baseline. For instance, with 1024 calibration samples, perplexities on WikiText-2 and C4 rise slightly from 7.19 and 11.17 to 7.51 and 11.80, respectively. These findings highlight FuseGPT’s robustness under both pruning and quantization, demonstrating that its block-level fusion strategy effectively preserves critical weights for downstream prediction even when weights are stored at a lower numerical precision.
5 Conclusion
In this paper, we presented FuseGPT, a structured pruning technique that recycles redundant transformer blocks to accelerate Generative Pre-trained Transformers (GPTs) without compromising quality. Guided by the novel Macro Inference (MI) metric, FuseGPT identifies less influential blocks and seamlessly merges them with neighboring ones, preserving crucial pre-trained knowledge. FuseGPT retains low perplexities across various datasets while reducing latency and memory costs, exceeding the competitiveness of prior structured and unstructured pruning solutions. By balancing performance with practical deployment needs, FuseGPT lays the groundwork for more scalable and efficient GPT-based models, opening up broader avenues for further research.